# Gautschi-type and implicit–explicit integrators for constrained wave equations

### v1

@@ 4.3. A Gautschi-type integration scheme
Based on the solution formula (4.9) we are now able to construct a Gautschi-type integrator. For this, we fix the nonlinear term in the integral of (4.9) by its value on the left limit, i.e., by $f(t_{n},u^{n})-\mathcal{B}^{-}\ddot{g}^{n}$. With
@@
We would like to emphasize that the Gautschi-type integrator (4.10) is a two-step scheme. Hence, we need to provide approximations $u^{0}$ and $u^{1}$ to apply the scheme. For $u^{0}$ we can directly insert $u(0)$. For the second value, we need to ensure $u^{1}=u(\tau)+\mathcal{O}(\tau^{3})$ in order to allow a second-order accuracy of the scheme, see Theorem 4.4 below. As an example, one may use a Taylor series approach in combination with (4.8), leading to
@@ 4.4. Convergence analysis
As preparation for the upcoming convergence proof, we summarize some facts on $\cos(t\Omega)$, $\operatorname{sinc}(t\Omega)\colon[0,\infty)\to\mathcal{L}(\mathcal{V})$. Recall that these time-dependent operators should be understood as part of the solution mapping, see (4.6) and (4.7).
@@
Note that we use $2kt=(2k+1)t+(-t)$ within the first application of Lemma 4.2b) in combination with an index shift. Moreover, we would like to emphasize at this point that $\sin(t\Omega)$ may not be defined. However, due to $x\operatorname{sinc}(x)=\sin(x)$, the above expressions only contain $\Omega^{2}$, which makes it well-defined. If, on the other hand, $n$ is odd, then we get
$X_{n}$ ${=}2\cos(t\Omega)X_{n-1}-X_{n-2}$
${=}2\cos(t\Omega)\Big{(}\operatorname{id}+2\sum_{k=1}^{(n-1)/2}\cos(2kt\Omega)\Big{)}-2\sum_{k=1}^{(n-1)/2}\cos\big{(}(2k-1)\,t\Omega\big{)}$
@@
${=}2\sum_{k=1}^{(n+1)/2}\cos\big{(}(2k-1)\,t\Omega\big{)}.\qed$
@@
Let Assumptions 2.1, 2.2, and 2.6 (a) be satisfied and assume consistent initial data $u^{0}\in\mathcal{V}$, $w^{0}\in\mathcal{H}$ leading to $u^{1}$ satisfying $\|u^{1}-u(t_{1})\|_{\mathcal{H}}\leq\tau^{3}E_{\text{init}}$. Further assume right-hand sides
@@
Considering (4.9) at times $t_{n+1}$ and $t_{n-1}$ (and omitting the restriction operator $\pi_{\text{ker}}$ for simplicity) yields
In combination with the approximate solution and formula (4.10), we obtain for the error $e^{n}\coloneqq u^{n}-u(t_{n})$ the equation
@@
with $L_{\mathcal{H}}$ denoting the Lipschitz constant of $f$ from Assumption 2.6. Also the integral term in (4.12) equals the solution of (4.8) evaluated at $t=\tau$ for homogeneous initial data but now with right-hand side
@@
An application of [GASSS98, Lem. 2] finally gives the stated result. ∎
@@ 4.5. Practical Implementation
Before presenting numerical examples, we need to discuss how the introduced Gautschi-type integrator (4.10) can be implemented. We first perform a spatial discretization, e.g., by finite elements, leading to matrices $A$ and $B$ as discrete versions of $\mathcal{A}$ and $\mathcal{B}$, respectively. Moreover, we obtain a mass matrix $M$, which appears in front of the second-order term.
@@
Finally, we are concerned with the computation of $\cos(\tau\Omega_{\text{ker}})x_{0}$ for a given vector $x_{0}\in\ker B\subseteq\mathbb{R}^{n}$. Since the Taylor series of $\cos(\tau\Omega_{\text{ker}})$ only contains even powers of $\Omega_{\text{ker}}$, we do not need to compute $\Omega_{\text{ker}}$ itself. It suffices to compute $A_{\text{ker}}$ or rather its action on a vector. Hence, we consider the Krylov subspace spanned by the vectors $A_{\text{ker}}^{k}x_{0}$ for $k\geq 0$. Following the general idea of Krylov subspace methods (see, e.g., [Saa92, EE06]), we introduce
@@
with unit basis vector $e_{1}\in\mathbb{R}^{r}$. For the computation of $\cos(\tau H_{r})$, which is of moderate size, one may use methods described in [Hig08, Ch. 12]. For our numerical examples, we use the algorithm introduced in [AMHR15].
@@
Using the Arnoldi method for the computation of the matrix cosine leaves open the question of stopping criteria. Here, one may apply the determination criterion presented in [HPS+15]. In the numerical experiment of Section 5, $r=3$ was already sufficient.
@@ 5. Numerical Example
Here, $\Delta_{\Gamma}$ denotes the Laplace–Beltrami operator, i.e., the Laplacian along the boundary; cf. [GT01, Ch. 16.1]. For the inhomogeneities, we set
@@
To show well-posedness of (5.1), the system can be written in an abstract form; see [HK20, HL20]. To obtain the PDAE structure (2.1), we consider an alternative formulation as introduced in [Alt19, Alt23]. For this, we introduce a second variable $p\coloneqq u|_{\Gamma}$, which is only defined on the boundary. With this, we can interpret (5.1a) as a dynamic equation for $u$ and (5.1b) as a dynamic equation for $p$, which are coupled through the mentioned connection. We further define
$\mathcal{V}\coloneqq H^{1}(\Omega)\times H^{1}(\Gamma),\qquad\mathcal{Q}\coloneqq H^{-1/2}(\Gamma).$
@@
which has exactly the structure of (2.1). Therein, the constraint operator is given by $\mathcal{B}\colon\mathcal{V}\to\mathcal{Q}^{*}=H^{1/2}(\Gamma)$, $\mathcal{B}(u,p)\coloneqq p-\operatorname{trace}u$ with the usual trace operator. The differential operators are given by $\mathcal{A}_{\Omega}\colon H^{1}(\Omega)\to[H^{1}(\Omega)]^{*}$ and $\mathcal{A}_{\Gamma}\colon H^{1}(\Gamma)\to[H^{1}(\Gamma)]^{*}$,
@@
Note that, in contrast to (5.1), system (5.2) contains two additional variables, namely $p$ for the trace values and a Lagrange multiplier $\lambda\colon[0,T]\to\mathcal{Q}$ which enforces the constraint.
@@
The spatial meshes are generated by DistMesh [PS04] and we use piecewise linear finite element approximations. Moreover, we set $T=1$ and, as initial condition,
@@
Accordingly, $p(0)$ and $\dot{p}(0)$ are defined in a consistent manner. Within the experiment, we have used six spatial level refinements, leading to a mesh size of around $h\approx 0.0672$. The reference solution is computed on the same spatial grid by the Gautschi-type integrator with Krylov subspace dimension $r=10$ and a time step size of $\tau_{\text{ref}}=2^{-13}$.
@@
We further observe the proven second-order accuracy of the IMEX scheme (3.2). For the Gautschi-type integrator (4.10), we present three versions which differ by the dimension of the Krylov subspaces used for the computation of the matrix cosine. For dimension $r=1$, the error stagnates but already for $r=2$ the errors show second-order convergence. The results improve once more for $r=3$ but remain unchanged for larger dimensions $r>3$. If one considers the $L^{\infty}(0,T;H^{1}(\Omega))$-norm, in which the IMEX Crank–Nicolson error was estimated in Theorem 3.1, we observe the same orders but with larger error constants.
$10^{-3}$$10^{-2}$$10^{-1}$$10^{-7}$$10^{-5}$$10^{-3}$$10^{-1}$time step size $\tau$error in $u$IMEX EulerIMEX CNGautschi ($r=1$)Gautschi ($r=2$)Gautschi ($r=3$)
Figure 5.1. Convergence history for different time stepping schemes. Plot shows error in $u$, measured in the $L^{\infty}(0,T;L^{2}(\Omega))$-norm. The gray lines indicate orders $1$ (dashed) and $2$ (dotted).

### v2

@@ 4.3. A Gautschi-type integration scheme
Considering the solution formula (4.9) at times $t_{n+1}$ and $t_{n-1}$ (and omitting the restriction operator $\pi_{\text{ker}}$ for simplicity), we get
Now, we fix the nonlinear term in the integral by its value on the left limit, i.e., by $f(t_{n},u^{n})-\mathcal{B}^{-}\ddot{g}^{n}$. With
@@
We would like to emphasize that the Gautschi-type integrator (4.10) is a two-step scheme. Hence, we need to provide approximations $u^{0}$ and $u^{1}$ to apply the scheme. For $u^{0}$ we can directly insert $u(0)$. For the second value, we need to ensure $u^{1}=u(\tau)+\mathcal{O}(\tau^{3})$ in order to allow a second-order accuracy of the scheme; see Theorem 4.4 below. In fact, due to the second order structure of the problem, in the error accumulation we indeed loose one order of $\tau$ on the initial value and two orders in the defect. As an example, one may use a Taylor series approach in combination with (4.8), leading to
@@ 4.4. Convergence analysis
As preparation for the upcoming convergence proof, we summarize some facts on $\cos(t\Omega)$, $\operatorname{sinc}(t\Omega)\colon[0,\infty)\to\mathcal{L}(\mathcal{V})$. Recall that these time-dependent operators should be understood as part of the solution mapping; see (4.6) and (4.7).
@@
Note that we use 2kt=(2k+1)t+(t)2kt=(2k+1)t+(-t) within the first application of Lemma 4.2b) in combination with an index shift. Moreover, we would like to emphasize at this point that $\sin(t\Omega)$ may not be defined. However, due to $x\operatorname{sinc}(x)=\sin(x)$, the above expressions only contain $\Omega^{2}$, which makes it well-defined. If, on the other hand, $n$ is odd, then we get
[…]
@@
Let Assumptions 2.1, 2.2, and 2.6 (a) be satisfied and assume consistent initial data $u^{0}\in\mathcal{V}$, $w^{0}\in\mathcal{H}$ leading to $u^{1}$ satisfying $\|u^{1}-u(t_{1})\|_{\mathcal{H}}\leq\tau^{3}E_{\text{init}}$. Further, assume right-hand sides
@@
As in the beginning of Section 4.3, we consider (4.9) at times $t_{n+1}$ and $t_{n-1}$. In combination with the approximate solution and formula (4.10), we obtain for the error $e^{n}\coloneqq u^{n}-u(t_{n})$ the equation
@@
with $L_{\mathcal{H}}$ denoting the Lipschitz constant of $f$ from Assumption 2.6. Also, the integral term in (4.12) equals the solution of (4.8) evaluated at $t=\tau$ for homogeneous initial data but now with right-hand side
@@
An application of [GarSS98, Lem. 2] finally gives the stated result. ∎
@@ 4.5. Practical Implementation
Before presenting numerical examples, we need to discuss how the introduced Gautschi-type integrator (4.10) can be implemented. For this, we assume a spatial discretization leading to the matrices $A$, $B$, and $M$; cf. Remark 3.3.
@@
Finally, we are concerned with the computation of $\cos(\tau\Omega_{\text{ker}})x_{0}$ for a given vector $x_{0}\in\ker B\subseteq\mathbb{R}^{n}$. Since the Taylor series of $\cos(\tau\Omega_{\text{ker}})$ only contains even powers of $\Omega_{\text{ker}}$, we do not need to compute $\Omega_{\text{ker}}$ itself. It suffices to compute $A_{\text{ker}}$ or rather its action on a vector. Hence, we consider the Krylov subspace spanned by the vectors $A_{\text{ker}}^{k}x_{0}$ for $k\geq 0$. Following the general idea of Krylov subspace methods (see, e.g., [Saa92, EieE06]), we introduce
@@
with unit basis vector $e_{1}\in\mathbb{R}^{r}$. For the computation of $\cos(\tau H_{r})$, which is of moderate size, one may use methods described in [Hig08, Ch. 12]. For our numerical examples, we use the algorithm introduced in [AlMHR15].
@@
Using the Arnoldi method for the computation of the matrix cosine leaves open the question of stopping criteria. Here, one may apply the determination criterion presented in [HocPSTW15]. In the numerical experiment of Section 5, $r=3$ was already sufficient.
@@ 5. Numerical Example
Here, $\Delta_{\Gamma}$ denotes the Laplace–Beltrami operator, i.e., the Laplacian along the boundary; cf. [GilT01, Ch. 16.1]. For the inhomogeneities, we set
@@
Note that $f_{\Omega}$ satisfies part (a) and (b) of Assumption 2.6, but $f_{\Gamma}$ only satisfies part (b).
To show well-posedness of (5.1), the system can be written in an abstract form; see [HipK20, HocL20]. To obtain the PDAE structure (2.1), we consider an alternative formulation as introduced in [Alt19, Alt23]. For this, we introduce a second variable $p\coloneqq u|_{\Gamma}$, which is only defined on the boundary. With this, we can interpret (5.1a) as a dynamic equation for $u$ and (5.1b) as a dynamic equation for $p$, both coupled through the mentioned connection. We further define
$\mathcal{V}\coloneqq H^{1}(\Omega)\times H^{1}(\Gamma),\qquad\mathcal{H}\coloneqq L^{2}(\Omega)\times L^{2}(\Gamma),\qquad\mathcal{Q}\coloneqq H^{-1/2}(\Gamma).$
@@
which has exactly the structure of (2.1). Therein, the constraint operator is given by $\mathcal{B}\colon\mathcal{V}\to\mathcal{Q}^{*}=H^{1/2}(\Gamma)$, $\mathcal{B}(u,p)\coloneqq p-\operatorname{trace}u$ with the usual trace operator. The kernel of $\mathcal{B}$ in $\mathcal{V}$ is given by
$\mathcal{V}_{\ker}=\ker\mathcal{B}=\big\{(u,p)\in\mathcal{V}\ |\ \operatorname{trace}u=p\big\},$
leading to $\mathcal{H}_{\ker}=\mathcal{H}$. Thus, we are in the modified setting of Remark 3.10. The differential operators are given by $\mathcal{A}_{\Omega}\colon H^{1}(\Omega)\to[H^{1}(\Omega)]^{*}$ and $\mathcal{A}_{\Gamma}\colon H^{1}(\Gamma)\to[H^{1}(\Gamma)]^{*}$,
@@
Note that, in contrast to (5.1), system (5.2) contains two additional variables, namely $p$ for the trace values and a Lagrange multiplier $\lambda\colon[0,T]\to\mathcal{Q}$, which enforces the constraint.
@@
The spatial meshes are generated by DistMesh [PerS04], and we use piecewise linear finite element approximations. Moreover, we set $T=1$ and, as initial condition,
@@
Accordingly, $p(0)$ and $\dot{p}(0)$ are defined consistently. Within the experiment, we have used six spatial level refinements, leading to a mesh size of around $h\approx 0.0672$. The reference solution is computed on the same spatial grid by the Gautschi-type integrator with Krylov subspace dimension $r=10$ and a time step size of $\tau_{\text{ref}}=2^{-13}$.
@@
We further observe the proven second-order accuracy of the IMEX scheme (3.2). For the Gautschi-type integrator (4.10), we present three versions that differ by the dimension of the Krylov subspaces used for the computation of the matrix cosine. For dimension $r=1$, the error stagnates but already for $r=2$ the errors show second-order convergence. The results improve once more for $r=3$ but remain unchanged for larger dimensions $r>3$. We note that the choice of $f_{\Gamma}$ is covered by the assumption for the IMEX scheme, but not the results for the Gautschi method. However, different choices lead to similar convergence plots. Finally, Figure 5.1 also shows the results for the fully explicit leapfrog scheme. Here, we observe the expected second-order convergence as soon as the scheme gets stable for sufficiently small time step sizes.
If one considers the $L^{\infty}(0,T;H^{1}(\Omega))$-norm, in which the IMEX Crank–Nicolson error was estimated in Theorem 3.1, we observe the same orders but with larger error constants.
[…]
Figure 5.1. Convergence history for different time stepping schemes. Plot shows error in $u$, measured in the $L^{\infty}(0,T;L^{2}(\Omega))$-norm. The gray lines indicate orders 11 (dashed) and 22 (dotted).